# Theory on Mixture-of-Experts in Continual Learning

### v2

@@ 5 Theoretical results on forgetting and generalization
$\textstyle F_{t}=\frac{1}{t-1}\sum_{\tau=1}^{t-1}(\mathcal{E}_{\tau}(\bm{w}_{t}^{(m_{\tau})})-\mathcal{E}_{\tau}(\bm{w}_{\tau}^{(m_{\tau})})).$ (15)
@@
$\textstyle G_{T}=\frac{1}{T}\sum_{\tau=1}^{T}\mathcal{E}_{\tau}(\bm{w}_{T}^{(m_{\tau})}).$ (16)
In the following, we present explicit forms of the above two metrics for learning with a single expert (i.e., $M=1$) as a benchmark [20]. Here we define $r:=1-\frac{s}{d}$ as the overparameterization ratio.
@@
Note that the setting here (with $M=1$) is slightly different from that in [20], because we have $N$ tasks in total. Hence a proof of Proposition 4 is provided in Appendix H. Proposition 4 implies that distinct tasks with large model gap $\sum_{n\neq n^{\prime}}\|\bm{w}_{n}-\bm{w}_{n}^{\prime}\|^{2}$ lead to poor performance of both $\mathbb{E}[F_{t}]$ in Eq. 17 and $\mathbb{E}[G_{t}]$ in Eq. 18.
@@ 5.1 Case I: More experts than tasks
1) Forgetting. If $t\leq T_{1}$, in Eq. 19, the coefficient $r^{L_{t}^{(m_{\tau})}}-r^{L_{\tau}^{(m_{\tau})}}$ of the term $\sum_{n=1}^{N}\|\bm{w}_{n}\|^{2}$ is smaller than 00 because $L_{t}^{(m_{\tau})}\geq L_{\tau}^{(m_{\tau})}$ and $r<1$, indicating that the training of new tasks enhances the performance of old tasks due to the repeated task arrivals in this phase. Meanwhile, the coefficient of model gap $\sum_{n\neq n^{\prime}}\|\bm{w}_{n^{\prime}}-\bm{w}_{n}\|^{2}$ is greater than 00, indicating that the forgetting is due to experts’ exploration of distinct tasks. However, as stated in Proposition 1, once the expert models converge at $t=T_{1}$, subsequent training on new task arrivals does not lead to forgetting of previous tasks. Therefore, for $t\in\{T_{1}+1,\cdots,T\}$, $\mathbb{E}[F_{t}]=\frac{T_{1}-1}{t-1}\mathbb{E}[F_{T_{1}}]$ decreases with $t$ and converges to zero as $T\rightarrow\infty$. This trend indicates that, compared to the oscillated forgetting in Eq. 17 of a single expert, the MoE model can significantly reduce expected forgetting for CL. Furthermore, a decrease in task similarity, i.e., larger model gaps, further amplifies the learning benefit of the MoE model.
@@ 6 Experiments
In this section, we conduct extensive experiments for both linear models and DNNs to support our theoretical analysis. Note that compared to most theoretical studies in CL (e.g., [19, 47]), our work has already included a more comprehensive experimental investigation. Due to the space limit, we delegate the experimental details and additional experiments on the impact of the termination threshold and load balancing to Appendix A.
Key design of early termination. In the first experiment, we aim to check the necessity of terminating the update of $\mathbf{\Theta}_{t}$ in Line 11 of Algorithm 1. Here we set $T=2000,N=7,K=3$ and vary expert number $M\in\{1,5,10,20\}$. As depicted in Figure 2(a) and Figure 2(c), both forgetting and generalization error first increase due to the expert exploration and then converge to almost zero for all MoE models with termination of update, verifying Theorem 1 and Theorem 2. In stark contrast, learning without termination leads to poor performance with large oscillations in Figure 2(b) and Figure 2(d), as the router selects the wrong expert for a new task arrival after the continual update of $\mathbf{\Theta}_{t}$. In addition, in both Figure 2(a) and Figure 2(c), the MoE model significantly improves the performance of CL compared to a single model. The comparison between $M=10$ and $M=20$ also indicates that adding extra experts delays the convergence if $M>N$, which does not improve learning performance, verifying our analysis in Theorem 1 and Theorem 2.
@@
Real-data validation. Finally, we extend our Algorithm 1 and insights from linear models to DNNs by conducting experiments on the MNIST dataset [48]. We set $N=3$ and vary $M\in\{1,4,7\}$. In each training round, we obtain the feature matrix by averaging $s=100$ training data samples. To diversify the model gaps of different tasks, we transform the $d\times d$ matrix into a $d\times d$ dimensional normalized vector to serve as input for the gating network. Then we calculate the variance $\sigma_{0}$ of each element across all tasks from the input vector, which is then used for parameter setting in Algorithm 1. Figure 3 illustrates that our theoretical insights from linear models also hold for DNNs, in terms of the impact of MoE and early termination on the performance of CL in practice.
@@
Figure 3: Learning performance under MNIST datasets [48]. Here we set $N=3$ and
$M\in\{1,4,7\}$
@@ 7 Conclusion
In this work, we conducted the first theoretical analysis of MoE and its impact on learning performance in CL, focusing on an overparameterized linear regression problem. We proved that the MoE model can diversify experts to specialize in different tasks, while its router can learn to select the right expert for each task and balance the loads across all experts. Then we demonstrated that, under CL, terminating the updating of gating network parameters after sufficient training rounds is necessary for system convergence. Furthermore, we provided explicit forms of the expected forgetting and overall generalization error to assess the impact of MoE. Finally, we conducted experiments on real datasets using DNNs to show that certain insights can extend beyond linear models.
@@ A.1 Experiments compute resources
Operating system: macOS Sonoma 14.2.1.
Type of CPU: 2.6 GHz 6-Core Intel Core i7
Type of GPU: Intel UHD Graphics 630 1536 MB
Memory: 16 GB 2667 MHz DDR4
Time of Execution: $40$ minutes for Figure 2, $100$ minutes for Figure 3, and $80$ minutes for Figure 4 and Figure 5.
@@ A.3 Experimental details of Figure 3
Datasets. We use the MNIST dataset, selecting $100$ samples randomly for training and $1000$ samples for testing at each training round.
@@
Task setups. We set ground truth pool $\mathcal{W}=\{(1),(4),(7)\}$ with $N=3$ tasks and $T=60$ training rounds. Before conducting the experiments in Figure 3, we first randomly generate the order of task arrivals $[n_{t}]_{t\in[T]}$, where $n_{t}$ at each round $t$ is randomly drawn from $\{1,2,3\}$ with equal probability $\frac{1}{3}$. Then, we conduct the two experiments (i.e., with and without termination) under the same order of task arrivals.
A.4 Experiments on termination threshold and load balance
@@ E.4 Proof of Lemma 10
$|\pi_{m}(\mathbf{X}_{t},\mathbf{\Theta}_{t})-\pi_{m^{\prime}}(\mathbf{X}_{t},\mathbf{\Theta}_{t})|$ $=\Big{|}\pi_{m^{\prime}}(\mathbf{X}_{t},\mathbf{\Theta}_{t})\exp{(h_{m}(\mathbf{X}_{t},\bm{\theta}_{t}^{(m)})-h_{m^{\prime}}(\mathbf{X}_{t},\bm{\theta}_{t}^{(m^{\prime})})}-\pi_{m^{\prime}}(\mathbf{X}_{t},\mathbf{\Theta}_{t})\Big{|}$
@@ E.6 Proof of Lemma 12
In the case of $M>N$, as $|\mathcal{M}_{n}|=1$, if task $n_{t}$ with $n_{t}=n$ is routed to the correct expert $m_{t}\in\mathcal{M}_{n}$, we have $\bm{w}^{m_{t}}_{t}=\bm{w}^{m_{t}}_{t-1}$ by (5). Consequently, the caused locality loss $\mathcal{L}^{loc}_{t}(\bm{\Theta}_{t},\mathcal{D}_{t})=0$, based on its definition in Eq. 6.
@@ E.7 Proof of Lemma 13
We first focus on the $M>N$ case to prove Lemma 13. Based on Eq. 35 in Proposition 1, we obtain $\pi_{m}(\mathbf{X}_{n},\mathbf{\Theta}_{T_{1}})-\pi_{m}(\mathbf{X}_{n^{\prime}},\mathbf{\Theta}_{T_{1}})=\Omega(\sigma_{0}^{0.5})$ for $m\in\mathcal{M}_{n}$, given $\|\bm{\theta}_{t}^{(m)}\|_{\infty}=\mathcal{O}(\sigma_{0}^{0.5})$ derived in Lemma 6. To prove Eq. 33, we will prove that $\pi_{m}(\mathbf{X}_{n},\mathbf{\Theta}_{t})-\pi_{m}(\mathbf{X}_{n^{\prime}},\mathbf{\Theta}_{t})=\Omega(\sigma_{0}^{0.5})$ holds for any $t\geq T_{1}+1$.
@@
$\|\bm{\theta}_{T_{2}}^{(m)}-\bm{\theta}_{T_{1}+1}^{(m)}\|_{\infty}$ $<\mathcal{O}(\sigma_{0}^{0.5})-\|\nabla_{\bm{\theta}_{T_{1}}^{(m)}}\mathcal{L}_{T_{1}}^{task}\|_{\infty}\cdot\eta\cdot(T_{2}-T_{1})$

### v3

@@ 5 Theoretical results on forgetting and generalization
$F_{t}=\frac{1}{t-1}\sum_{\tau=1}^{t-1}(\mathcal{E}_{\tau}(\bm{w}_{t}^{(m_{\tau})})-\mathcal{E}_{\tau}(\bm{w}_{\tau}^{(m_{\tau})})).$ (15)
@@
$G_{T}=\frac{1}{T}\sum_{\tau=1}^{T}\mathcal{E}_{\tau}(\bm{w}_{T}^{(m_{\tau})}).$ (16)
In the following, we present explicit forms of the above two metrics for learning with a single expert (i.e., $M=1$) as a benchmark (cf. [34]). Here we define $r:=1-\frac{s}{d}$ as the overparameterization ratio.
@@
Note that the setting here (with $M=1$) differs slightly from [34] as we have $N$ tasks in total. Hence a proof of Proposition 4 is provided in Appendix H. Proposition 4 implies that distinct tasks with large model gap $\sum_{n\neq n^{\prime}}\|\bm{w}_{n}-\bm{w}_{n}^{\prime}\|^{2}$ lead to poor performance of both $\mathbb{E}[F_{t}]$ in Eq. 17 and $\mathbb{E}[G_{t}]$ in Eq. 18, which is missing in the existing CL literature (e.g., [30]).
@@ 5.1 Case I: More experts than tasks
1) Forgetting. If $t\leq T_{1}$, in Eq. 19, the coefficient $r^{L_{t}^{(m_{\tau})}}-r^{L_{\tau}^{(m_{\tau})}}$ of the term $\sum_{n=1}^{N}\|\bm{w}_{n}\|^{2}$ is smaller than 00 because $L_{t}^{(m_{\tau})}\geq L_{\tau}^{(m_{\tau})}$ and $r<1$, indicating that the training of new tasks enhances the performance of old tasks due to the repeated task arrivals in this phase. Meanwhile, the coefficient of model gap $\sum_{n\neq n^{\prime}}\|\bm{w}_{n^{\prime}}-\bm{w}_{n}\|^{2}$ is greater than 00, indicating that the forgetting is due to experts’ exploration of distinct tasks. However, as stated in Proposition 1, once the expert models converge at $t=T_{1}$, training on newly arriving tasks with correct routing no longer causes forgetting of previous tasks. Consequently, for $t\in\{T_{1}+1,\cdots,T\}$, $\mathbb{E}[F_{t}]=\frac{T_{1}-1}{t-1}\mathbb{E}[F_{T_{1}}]$ decreases with $t$ and converges to zero as $T\rightarrow\infty$. This result highlights that, unlike the oscillatory forgetting observed in Eq. 17 for a single expert, the MoE model effectively minimizes expected forgetting in CL through its correct routing mechanism. Furthermore, a decrease in task similarity, i.e., larger model gaps, further amplifies the learning benefit of the MoE model.
@@ 6 Experiments
In this section, we present extensive experiments on both linear models and DNNs to validate our theoretical analysis. Due to space constraints, we include detailed experimental setups and additional results on datasets such as MNIST [27], CIFAR-100 [25] and Tiny ImageNet [26] in Appendix A.
Key design of early termination. In the first experiment, we aim to check the necessity of terminating the update of $\mathbf{\Theta}_{t}$ in Line 11 of Algorithm 1. Here we set $T=2000,N=6,K=3$ and vary expert number $M\in\{1,5,10,20\}$. As depicted in Figure 2(a) and Figure 2(c), both forgetting and generalization error first increase due to the expert exploration and then converge to almost zero for all MoE models with termination of update, verifying Theorem 1 and Theorem 2. In stark contrast, learning without termination leads to poor performance with large oscillations in Figure 2(b) and Figure 2(d), as the router selects the wrong expert for a new task arrival after the continual update of $\mathbf{\Theta}_{t}$. In addition, in both Figure 2(a) and Figure 2(c), the MoE model significantly improves the performance of CL compared to a single model. The comparison between $M=10$ and $M=20$ also indicates that adding extra experts delays the convergence if $M>N$, which does not improve learning performance, verifying our analysis in Theorem 1 and Theorem 2.
@@
Real-data validation. Finally, we extend our Algorithm 1 and insights from linear models to DNNs by conducting experiments on the CIFAR-10 dataset ([25]). The details of our experiment setup is given in Section A.3. We set $K=4,N=300$ and vary $M\in\{1,4,12\}$. In each training round, to diversify the model gaps of different tasks, we transform the $d\times d$ matrix into a $d\times d$ dimensional normalized vector to serve as input for the gating network. Then we calculate the variance $\sigma_{0}$ of each element across all tasks from the input vector, which is then used for parameter setting in Algorithm 1. Figure 3 illustrates that our theoretical insights from linear models also hold for DNNs, in terms of the impact of MoE and early termination on the performance of CL in practice.
@@
Figure 3: The dynamics of overall generalization error and test accuracy under the CIFAR-10 dataset ([25]). Here we set $K=4,N=300$ and $M\in\{1,4,12\}$.
@@ 7 Conclusion
In this work, we conducted the first theoretical analysis of MoE and its impact on learning performance in CL, focusing on an overparameterized linear regression problem. We establish the benefit of MoE over a single expert by proving that the MoE model can diversify its experts to specialize in different tasks, while its router learns to select the right expert for each task and balance the loads across all experts. Then we demonstrated that, under CL, terminating the updating of gating network parameters after sufficient training rounds is necessary for system convergence. Furthermore, we provided explicit forms of the expected forgetting and overall generalization error to assess the impact of MoE. Interestingly, adding more experts requires additional rounds before convergence, which may not enhance the learning performance. Finally, we conducted experiments on real datasets using DNNs to show that certain insights can extend beyond linear models.
Acknowledgments
This work has been supported in part by the U.S. National Science Foundation under the grants: NSF AI Institute (AI-EDGE) 2112471, CNS-2312836, and was sponsored by the Army Research Laboratory under Cooperative Agreement Number W911NF-23-2-0225. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Laboratory or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
This work was also supported in part by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 Grant with Award no. MOE-T2EP20121-0001; in part by SUTD Kickstarter Initiative (SKI) Grant with no. SKI 2021_04_07; and in part by the Joint SMU-SUTD Grant with no. 22-LKCSB-SMU-053.
The work of Yingbin Liang was supported in part by the U.S. National Science Foundation with the grants RINGS-2148253, CNS-2112471, and ECCS-2413528.
We sincerely thank Qinhang Wu for his invaluable assistance in conducting extensive experiments during our rebuttal, which greatly contributed to the acceptance of this paper.
@@ A.1 Experiments compute resources
Operating system: Red Hat Enterprise Linux Server 7.9 (Maipo)
Type of CPU: 2.9 GHz 48-Core Intel Xeon 8268s
Type of GPU: NVIDIA Volta V100 w/32 GB GPU memory
@@ A.3 Experimental details of Figure 3
Datasets. We use the CIFAR-10 ([25]) dataset, selecting $512$ samples randomly for training and $2000$ samples for testing at each training round.
DNN architecture and training details. We employ a non-pretrained ResNet-18 as our base model. Each task is learned using Adam with a learning rate governed by a cosine annealing schedule for 5 epoches, with a minibatch size of $32$, a weight decay of $0.005$. The initial learning rate is set to $0.0005$, and it is reduced to a minimum value of $10^{-6}$ over a total of $300$ rounds.
Task setups. We define the ground truth pool as $\mathcal{W}=\{(0),(4),(5),(9)\}$, representing $K=4$ clusters of tasks for recognizing the image classes airplane, deer, dog, truck, respectively. The experiment spans $T=300$ training rounds with $N=300$ tasks. We randomly generate the task arrival sequence $[n_{t}]_{t\in[T]}$, where each $n_{t}$ is drawn from ${(0),(4),(5),(9)}$ with equal probability $\frac{1}{4}$. We then conduct two experiments (with and without termination) using the same task arrival order. For each task $t\in[T]$, we randomly select its type (e.g., task $(0)$ for recognizing the airplane class image) and $512$ corresponding samples, ensuring that tasks have distinct distributions and features.
(a) With termination
(b) Without termination
Figure 4: The dynamics of forgetting under the CIFAR-10 dataset. Here we set $N=4$ and $M\in\{1,4\}$.
A.4 Experiments on the MNIST dataset
Datasets. We use the MNIST dataset [27], selecting $100$ samples randomly for training and $1000$ samples for testing at each training round.
@@ A.4 Experiments on the MNIST dataset
Task setups. We define the ground truth pool as $\mathcal{W}=\{(1),(4),(7)\}$, representing $K=3$ clusters of tasks for recognizing the numbers 1, 4, and 7, respectively. The experiment spans $T=60$ training rounds with $N=60$ tasks. Before the experiments in Figure 5, we randomly generate the task arrival sequence $[n_{t}]_{t\in[T]}$, where each $n_{t}$ is drawn from ${(1),(4),(7)}$ with equal probability $\frac{1}{3}$. We then conduct two experiments (with and without termination) using the same task arrival order. For each task $t\in[T]$, we randomly select its type (e.g., task $(1)$ for recognizing the number 1) and 100 corresponding samples, ensuring that tasks have distinct distributions and features.
(a) With termination.
(b) Without termination.
(c) With termination.
(d) Without termination.
Figure 5: Learning performance under the MNIST dataset ([27]). Here we set $K=3,N=60$ and
$M\in\{1,4,7\}$
A.5 Experiments on the CIFAR-100 dataset
Datasets. We use the CIFAR-100 ([25]) dataset, selecting $192$ samples randomly for training and $600$ samples for testing at each training round.
DNN architecture and training details. They are the same as the experiments on the CIFAR-10 dataset in Section A.3.
Task setups. We define the ground truth pool as $\mathcal{W}=\{(28),(40),(52),(72),(79),(99)\}$, representing $K=6$ clusters of tasks for recognizing the image classes telephone, bee, mountain, bear, turtle, tractor respectively. The experiment spans $T=350$ training rounds with $N=350$ tasks. We randomly generate the task arrival sequence $[n_{t}]_{t\in[T]}$, where each $n_{t}$ is drawn from ${(28),(40),(52),(72),(79),(99)}$ with equal probability $\frac{1}{6}$. We then conduct two experiments (with and without termination) using the same task arrival order. For each task $t\in[T]$, we randomly select its type (e.g., task $(28)$ for recognizing the telephone class image) and $192$ corresponding samples, ensuring that tasks have distinct distributions and features.
(a) With termination
(b) Without termination
Figure 6: Learning performance under the CIFAR-100 dataset. Here we set $N=6$ and $M\in\{1,8,10\}$.
Table 1: The average incremental accuracy under the CIFAR-100 dataset.
Expert number Test accuracy (%)
$M=1$ $17.1$
$M=8$ $25.5$
$M=8$ w/o ET $10.3$
$M=10$ $32.9$
$M=10$ w/o ET $9.1$
A.6 Experiments on the Tiny ImageNet dataset
Datasets. We use the Tiny ImageNet ([26]) dataset, selecting $192$ samples randomly for training and $300$ samples for testing at each training round.
DNN architecture and training details. They are the same as the experiments on the CIFAR-10 dataset in Section A.3.
Task setups. We define the ground truth pool as $\mathcal{W}=\{(20),(50),(83),(145),(168),(179)\}$, representing $K=6$ clusters of tasks for recognizing six disjoint image classes from Tiny Imagenet respectively. The experiment spans $T=300$ training rounds with $N=300$ tasks. We randomly generate the task arrival sequence $[n_{t}]_{t\in[T]}$, where each $n_{t}$ is drawn from ${(20),(50),(83),(145),(168),(179)}$ with equal probability $\frac{1}{6}$. We then conduct two experiments (with and without termination) using the same task arrival order. For each task $t\in[T]$, we randomly select its type (e.g., task $(20)$) and $192$ corresponding samples, ensuring that tasks have distinct distributions and features.
(a) With termination
(b) Without termination
Figure 7: Learning performance under the Tiny ImageNet dataset. Here we set $N=6$ and $M\in\{1,8,10\}$.
Table 2: The average incremental accuracy under the Tiny ImageNet dataset.
Expert number Test accuracy (%)
$M=1$ $17.4$
$M=8$ $37.6$
$M=8$ w/o ET $10.5$
$M=10$ $28.3$
$M=10$ w/o ET $10.3$
A.7 Experiments on termination threshold and load balance
@@ E.4 Proof of Lemma 10
$|\pi_{m}(\mathbf{X}_{t},\mathbf{\Theta}_{t})-\pi_{m^{\prime}}(\mathbf{X}_{t},\mathbf{\Theta}_{t})|$ $=\Big{|}\pi_{m^{\prime}}(\mathbf{X}_{t},\mathbf{\Theta}_{t})\exp{(h_{m}(\mathbf{X}_{t},\bm{\theta}_{t}^{(m)})-h_{m^{\prime}}(\mathbf{X}_{t},\bm{\theta}_{t}^{(m^{\prime})}))}-\pi_{m^{\prime}}(\mathbf{X}_{t},\mathbf{\Theta}_{t})\Big{|}$
@@ E.6 Proof of Lemma 12
In the case of $M>N$, as $|\mathcal{M}_{n}|=1$, if task $n_{t}$ with $n_{t}=n$ is routed to the correct expert $m_{t}\in\mathcal{M}_{n}$, we have $\bm{w}^{m_{t}}_{t}=\bm{w}^{m_{t}}_{t-1}$ by Eq. 5. Consequently, the caused locality loss $\mathcal{L}^{loc}_{t}(\bm{\Theta}_{t},\mathcal{D}_{t})=0$, based on its definition in Eq. 6.
@@ E.7 Proof of Lemma 13
We first focus on the $M>N$ case to prove Lemma 13. Based on Eq. 35, we obtain $\pi_{m}(\mathbf{X}_{n},\mathbf{\Theta}_{T_{1}})-\pi_{m}(\mathbf{X}_{n^{\prime}},\mathbf{\Theta}_{T_{1}})=\Omega(\sigma_{0}^{0.5})$ for $m\in\mathcal{M}_{n}$, given $\|\bm{\theta}_{t}^{(m)}\|_{\infty}=\mathcal{O}(\sigma_{0}^{0.5})$ derived in Lemma 6. To prove Eq. 33, we will prove that $\pi_{m}(\mathbf{X}_{n},\mathbf{\Theta}_{t})-\pi_{m}(\mathbf{X}_{n^{\prime}},\mathbf{\Theta}_{t})=\Omega(\sigma_{0}^{0.5})$ holds for any $t\geq T_{1}+1$.
@@
$\|\bm{\theta}_{T_{2}}^{(m)}\|_{\infty}$ $<\mathcal{O}(\sigma_{0}^{0.5})-\|\nabla_{\bm{\theta}_{T_{1}}^{(m)}}\mathcal{L}_{T_{1}}^{task}\|_{\infty}\cdot\eta\cdot(T_{2}-T_{1})$